OpenStudy (anonymous):

Write the equation of the line that is tangent to the circle x2 + y2 = 169 at the point (-5, 12).

OpenStudy (gebooors):

Where is center of the circle? Distance of point (-5,12) is one radius from center. Draw this radius and Make an equation for the line, y = km + b Tangent is perpendicular to the line, so use K1 * k2 = -1.

OpenStudy (anonymous):

I got -5/12x+169

OpenStudy (anonymous):

Wait 5/12x

OpenStudy (gebooors):

For radius k = -5/12. It is descending line. K for tangent should be positive

OpenStudy (anonymous):

No it's 5/12 It doesn't have a negative

OpenStudy (gebooors):

-5/12 * k =-1 K for tangent is 12/5, right?

OpenStudy (anonymous):

I really don't know all I know is that it doesn't have a negative

OpenStudy (gebooors):

K = -1 /(-5:12) = 12/5

OpenStudy (anonymous):

Is it 12/5x+12/169 ?

OpenStudy (gebooors):

Then use y -y0 = k(x- x0) Remember multiply both : kx and k* (-12)

OpenStudy (gebooors):

(12/5) x is right

OpenStudy (gebooors):

y - (-5) = 12/5 (x -12) Remember -5 in another side

OpenStudy (anonymous):

y=12/5x+12/169

OpenStudy (anonymous):

In standard form

OpenStudy (gebooors):

Right side is 12/5 x - 12*12 / 5 +5 Last one 5 comes from left side.

OpenStudy (anonymous):

Can you conform my answer?

OpenStudy (gebooors):

I got y = (12/5) x - 23 4/5 12* 12 = 144 -144/ 5 = -28 4/5, then add 5

OpenStudy (anonymous):

I got it wrong.

OpenStudy (gebooors):

Ok, I think you find the idea. I suggest you make a good Sketch of this. Good luck for your studies!

OpenStudy (ranga):

The tangent line is: y = 5/12x + b Point (-5, 12) is on that tangent line. 12 = 5/12 * (-5) + b 12 = -25/12 + b b = 12 + 25/12 = (144 + 25)/12 = 169/12 y = 5/12x + 169/12

OpenStudy (gebooors):

I got a sign error, when I moved 5 from left to right, Sorry!

OpenStudy (anonymous):

A plot is attached.